# FaithUn: Toward Faithful Forgetting in Language Models by Investigating the Interconnectedness of Knowledge

### v1

@@ 5.3 KLUE Mitigates Superficial Unlearning
We investigate superficial unlearning on all baselines with Gemma-2 (2B) in the FaithUn setting, as shown in Table 2. First, the default Gemma-2 model can correctly answer most questions, validating that FaithUn is well constructed. After the unlearning process, all baselines reach UA0.33absent0.33$\leq$ 0.33≤ 0.33, which validates that all methods can unlearn target knowledge. However, they fail to reliably remove implicit and interconnected knowledge, suggesting that their unlearning process is superficial. However, our method mitigates superficial unlearning and achieves faithful unlearning compared to other baselines, without significantly damaging the other knowledge to maintain (i.e., TA, SA, and MA). These results demonstrate that our method accurately identifies neurons relevant to contextual knowledge and successfully erases this knowledge. In addition, experiments on Gemma-2 (9B) and Llama-3.2 (3B) reveal that our method outperforms baselines, with results presented in Appendix B.4. We also conduct ablation studies for KLUE and demonstrate the validity of our proposed methods, as detailed in Appendix B.7.
@@
We demonstrate how the unlearning process affects other knowledge by plotting all scores from the Gemma-2 (2B) unlearning process against UA. As the UA score represents the progress of unlearning target knowledge (decreasing with unlearning), we can observe each method’s impact on other knowledge in Figure 2. All methods’ impact on the paraphrased questions (UA) shows a strong correlation with the UA score, suggesting that all methods pose robustness in dealing with different lexical forms (but hold the same meaning) of the questions. However, the baselines struggle to maintain other knowledge (TA and SA) and to forget interconnected knowledge (MA). In contrast, KLUE demonstrates robust unlearning performance by effectively forgetting interconnected knowledge and preserving other knowledge.
@@
5.5 The Impact of Neuron Localization
We adopt varying ratios of neuron selection $p\in\{0.01,0.05,0.1\}$ to investigate the effect of the knowledge neuron on Gemma-2 (2B). Also, we conduct experiments for the random neuron selection (i.e., $p\in\{0.01,0.05\}$). As a result, we reveal that a neuron ratio of 0.05 or 0.1 contributes to achieving faithful unlearning, showing that random neuron selection more significantly triggers superficial unlearning.
Figure 3: The ratio of neuron localization. We plot the accuracy of each metric for varying ratios of neuron localization.
@@ 5.6 Qualitative Analysis
We conduct a qualitative analysis for KLUE and GAret on Gemma-2 (2B). Both KLUE and GAret successfully unlearn the paraphrased question ($\mathcal{C}_{p}$), degrading label logits to 0.33. However, GAret has difficulty in unlearning the multi-hop question ($\mathcal{C}_{m}$), while mistakenly unlearns the same-answer questions ($\mathcal{C}_{s}$). On the other hand, KLUE faithfully unlearns them, mitigating superficial unlearning.
@@ 6 Conclusion
Our research identifies the limitation of existing unlearning benchmarks, which have not explored the interconnectedness of knowledge. To overcome this issue, we define superficial unlearning and propose a new benchmark, FaithUn, for evaluating generalization, multi-hop knowledge unlearning, and shortcut unlearning. Using this benchmark, we empirically demonstrate that existing unlearning methods are vulnerable to superficial unlearning. Furthermore, we propose a novel knowledge-localized unlearning method, KLUE, and demonstrate that it outperforms existing unlearning methods, effectively mitigating superficial unlearning. Our paper first illuminates the phenomenon of superficial unlearning and raises a new research question for a deeper analysis of the unlearning field.
@@
FaithUn is constructed based on Wikidata and is designed to investigate the unlearning of knowledge about famous people for application in various language models. Although knowledge is more interconnected for well-known individuals, our benchmark does not examine a broader range of people. Additionally, our study focuses solely on erasing the target label, leaving the issue of hallucinations in the unlearning process as future work, in line with prior studies.
@@ A.1 Detailed Dataset Comparison
In this section, we propose detailed comparisons with existing datasets to show the novelty of our benchmark clearly. Our benchmark aims to unlearn the knowledge of famous real-world entities, which can be prevalent in various language models, to consider the most practical situation of knowledge unlearning. Furthermore, our benchmark deals with the complex and interconnected nature of world knowledge; thus, we introduce three types of unlearning evaluation aspects (Generalization, Multi-hop knowledge unlearning, and Shortcut unlearning) for more deep analysis of real-world knowledge unlearning.
In summary, MUSE, KnowUnDo, and TOFU require fine-tuning to inject knowledge before unlearning, which may reduce their practicality. Additionally, only RWKU and our benchmark address real-world entities as targets for unlearning. Furthermore, most existing benchmarks, except for RWKU and our benchmark, have not considered related knowledge. However, RWKU has not explored the interconnections between knowledge and shortcut unlearning problems, which have become increasingly significant in unlearning due to its reliance on a limited number of training instances. For example, RWKU includes a target text for unlearning: ’Please forget Stephen King, who is an American author, renowned as the "King of Horror"’. It also contains a related knowledge question: "’Who plays the character Jack Torrance in the film The Shining?’". While the two questions are somewhat related, they represent independent pieces of knowledge, as they are not interconnected like multi-hop questions. In conclusion, the main contribution of our benchmark lies in evaluating whether unlearning methods perform faithful unlearning while considering knowledge interconnectedness within the real-world entity unlearning setting.
A.2 Dataset Format
Our FaithUn benchmark includes four types of datasets: the Base QA dataset ($\mathcal{C}$), the Paraphrased QA dataset ($\mathcal{C}_{p}$), the Multi-hop QA dataset ($\mathcal{C}_{m}$), and the Same-answer QA dataset ($\mathcal{C}_{s}$). Each instance in the Base QA dataset is matched with instances in other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) to examine the impact of unlearning on these datasets. Dataset statistics for the FaithUn benchmark are shown in Table 4. Examples in the FaithUn benchmark are shown in Table 10.
@@ A.2 Dataset Format
Table 4: Dataset statistics. FaithUn includes questions to be forgotten, collectively referred to as the Base QA dataset. Each question in this dataset forms a cluster, and questions from other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) are matched with those in the Base QA dataset, thereby being assigned to the corresponding cluster.
A.3 Details in Dataset Construction
A.3.1 Selected Entities and Relations.
A.3.2 Dataset Analysis.
@@ A.3 Details in Dataset Construction
A.3.3 Question Generation Prompt Templates
@@ B.3 More Details for the Experiments
We train and evaluate KLUE and other baselines on NVIDIA A100 GPU. For a fair comparison, we early stop the training procedure when UA0.33absent0.33$\leq$ 0.33≤ 0.33 is satisfied (random sampling from three answer options) to select the optimal model. Since a language model forgets all the knowledge when a learning rate is set too high, we have searched for the lowest learning rates, which can reach UA0.33absent0.33$\leq$ 0.33≤ 0.33 within the range $\lambda\in$ [1e-07, 3e-03]. We adopt batch size $\beta=4$ for all unlearning methods. We compute the final loss by weighted-summing the loss of forget samples and retaining samples. Specifically, we use $0.7$ and $1.0$ for the loss of forget samples and the retaining samples, respectively. We select $e=150$ as the maximum number of epochs in the training process.
@@
(1) Gradient Ascent (GA): Unlike the gradient descent used during the pre-training phase, GA (Jang et al., 2023; Yao et al., 2023) maximize the negative log-likelihood loss on the forget set. This method helps shift the model away from its original predictions, aiding in the unlearning process. (2) Gradient Ascent with a Retaining Loss (GAret): GA tends to unlearn other unrelated knowledge since it just maximizes the negative log-likelihood loss on the forget set. Therefore, we add an auxiliary retention loss to maximize the log-likelihood of the retaining set, securing the retention of other irrelevant knowledge. (3) Direct Preference Optimization (DPO): We adopt preference optimization to unlearn a language model to generate another answer. DPO (Rafailov et al., 2024; Jin et al., 2024) utilizes positive and negative instances to train the model. Therefore, we select the correct answer as the negative instance and also define two types of DPO methods to determine positive ones: (1) DPOmis (DPO using a mismatched answer) and (2) DPOneg (DPO using a rejection answer). DPOmis utilizes a randomly sampled answer as the positive instance. On the other hand, DPOrej utilizes a rejection text “I can’t answer the question." as the positive instance. Two DPO methods both aim to increase the probability of the positive instance compared to the negative one for the forget set, and they switch the positive and negative instances for training the retaining set. We search for $\beta_{DPO}\in[0.1,0.5]$ to optimize models. (4) NPO: NPO is a modified version of DPO that exclusively retains negative examples without positive ones. NPO can also be explained as a straightforward modification of the GA loss. We implement NPO (Zhang et al., 2024) for extended experiments. We search for $\beta_{NPO}\in[0.1,0.5]$ to optimize models. (5) RMU: We implement RMU (Li et al., 2024a), the representation learning-based unlearning model. For RMU experiments, we search for $\alpha_{RMU}\in\{20,50,100,150,200,300\}$ and use hyper-parameters $c=20$ and $l=7$, following the implementation details on the original GitHub Page [^4] https://github.com/centerforaisafety/wmdp. (6) Knowledge-Localized Unlearning (KLUE): We select only 5% of neurons from Feed-forward networks for the knowledge neuron localization, and update them using general gradient ascent with retention loss. We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term. The experiments analyzing varying hyper-parameters are shown in Section 5.5, Appendix B.6.2, and Appendix B.6.3.
B.4 KLUE successfully Mitigates Superficial Unlearning on Various Language Models
We conduct experiments on Llama-3.2 (3B) and Gemma-2 (9B) to reveal that our method is model-agnostic and generalizable. Table 5 shows the experimental results for Llama-3.2 (3B) and Gemma-2 (9B), respectively. In the Llama-3.2 (3B) results, all baselines are significantly exposed to superficial unlearning, while our method successfully outperforms other baselines in most metrics. Likewise, in the Gemma-2 (9B) results, the enhancement of mitigating superficial unlearning is dramatic after applying knowledge localization. These results demonstrate that our method can be applied to various language models and successfully mitigates superficial unlearning.
Llama-3.2 (3B) Default 90.91 87.28 85.65 50.57 -
GA 35.35 54.52 39.19 52.45 52.70
GAret 48.14 68.24 57.71 53.94 57.94
DPOrej 46.80 69.68 55.86 54.02 58.19
DPOmis 36.02 64.87 43.21 51.56 55.91
KLUE 45.79 77.58 65.12 53.99 62.73
Gemma-2 (9B) Default 91.92 89.87 86.57 48.07 -
GA 29.29 40.52 30.56 50.46 48.06
GAret 45.45 83.84 68.52 50.72 64.40
DPOrej 41.41 75.32 59.72 47.02 60.16
DPOmis 36.36 63.15 43.06 55.45 56.32
KLUE 40.40 89.83 81.48 60.48 72.85
Table 5: Llama-3.2 (3B) and Gemma-2 (9B) experimental results. We report the results of four metrics after unlearning the forget set (5%) in our settings. Bolded results indicate the best performance.
B.5 KLUE is Robust to Various Forget Sample Sizes
We conduct experiments on Gemma-2 (2B) for the varying sizes (i.e., 1%, 5%, and 10%) of the forget set to analyze the effect of unlearning samples. The experimental results are shown in Table 2 (5%) and Table 6 (1% and 10%). Our experiments reveal that existing methods undergo more problems in unlearning when the number of forget samples increases. Increasing the number of samples to be forgotten is more challenging since it requires modifying a greater amount of knowledge. However, our proposed method consistently outperforms other baselines; thus, the performance gap between our method and the baselines widens as the number of forget samples increases.
1% Default 72.22 85.34 71.43 54.18 -
GA 44.44 77.80 57.14 49.43 59.98
GAret 34.33 85.78 59.52 58.38 67.33
DPOrej 44.44 72.84 54.76 51.79 58.73
KLUE 36.11 85.34 63.09 59.77 68.02
5% Default 81.82 85.99 79.63 48.67 -
GA 36.02 48.92 37.19 48.34 49.61
GAret 34.01 77.58 66.51 53.21 65.82
DPOrej 41.75 68.96 63.58 49.67 60.11
KLUE 36.70 82.97 74.69 58.16 69.78
10% Default 83.84 85.34 76.82 50.05 -
GA 38.38 28.02 31.13 50.41 42.79
GAret 40.40 62.50 65.12 54.21 60.35
DPOrej 34.85 45.26 42.38 51.29 51.02
KLUE 40.91 81.03 69.98 59.18 67.32
Table 6: Unlearning experiments for varying forget sample sizes. We report the unlearning results for the varying number of forget set (i.g., 1% and 10%). The results for 5% are also found in Table 2.
B.6 Hyper-parameter Experiments
B.6.1 Sequential vs. Batch Unlearning
@@ B.6 Hyper-parameter Experiments
B.6.2 Hyper-parameter ($\alpha$) Experiments
@@
B.6.3 Neuron Ratio ($p$) Experiments
@@
B.6.4 The Various Prompt Templates Experiments
@@
B.7 Ablation Studies
We perform ablation experiments on each KLUE method using Gemma-2 2B to better understand their relative importance, as shown in Table 9. Regularization means the strategy of using the auxiliary regularization term for quantifying the knowledge relevance of each neuron, mitigating superficial unlearning. Localization corresponds to the entire knowledge neuron localization strategy. Sample Selection is the strategy that selects unforgotten samples by evaluating the memorization of each sample. For the ablation study, we remove each of them and measure the accuracy. As a result, we reveal that three methods significantly affect the faithfulness of unlearning. Regularization and Localization are useful to enhance SA and MA, mitigating superficial unlearning related to interconnected knowledge. These results demonstrate that selecting proper knowledge neurons to be updated is useful for handling interconnected knowledge. In addition, we illuminate that Sample Selection significantly increases TA and SA, mitigating overfitting and shortcut unlearning issues.

### v2

@@ 5.2 KLUE Mitigates Superficial Unlearning
Main experiments.
[…]
Forget ratios experiments.
We conduct experiments on Gemma-2 (2B) for the varying sizes (i.e., 1%, 5%, and 10%) of the forget set to analyze the effect of unlearning samples as shown in Table 4. The experiments reveal that existing methods encounter more problems in unlearning when the number of forgetting samples increases, since it requires modifying a greater amount of knowledge. However, our proposed method consistently outperforms other baselines; thus, the performance gap between our method and the baselines widens as the number of forget samples increases.
@@
We demonstrate how the unlearning process affects other knowledge by plotting all scores from the Gemma-2 (2B) unlearning process against UA. As the UA score represents the progress of unlearning target knowledge (decreasing with unlearning), we can observe each method’s impact on other knowledge in Figure 2. All methods’ impact on the paraphrased questions (UA) shows a strong correlation with the UA score, suggesting that all methods pose robustness in dealing with different lexical forms (but hold the same meaning) of the questions. However, the baselines struggle to maintain other knowledge (TA and SA) and to forget interconnected knowledge (MA). In contrast, KLUE demonstrates robust performance by effectively forgetting interconnected knowledge and preserving other knowledge.
5.4 The Impact of Neuron Localization
We adopt varying ratios of neuron selection $p\in\{0.01,0.05,0.1\}$ to examine the effect of the knowledge neuron on Gemma-2 (2B), shown in Figure 3. Also, we conduct experiments for the random neuron selection (i.e., $p\in\{0.01,0.05\}$). We show that KLUE achieves faithful unlearning with a neuron ratio of 0.05 or 0.1. In contrast, random neuron selection significantly shows superficial unlearning. This result reveals that the appropriate selection of knowledge neurons for unlearning is crucial to ensure the generalization of unlearning and the preservation of other knowledge.
@@ 5.4 The Impact of Neuron Localization
Figure 3: The ratio of neuron localization. We plot the accuracy of each metric for varying ratios of neurons.
@@ 5.5 Qualitative Analysis
We conduct a qualitative analysis for KLUE and GAret on Gemma-2 (2B), shown in Table 5. Both KLUE and GAret successfully unlearn the paraphrased question, degrading label probability to 0.33 (random guess). However, GAret has difficulty in multi-hop unlearning and mistakenly unlearns the same-answer questions. KLUE faithfully unlearns them, mitigating superficial unlearning.
5.6 Ablation Studies
We perform ablation studies on each KLUE method using Gemma-2 (2B) to better understand their relative importance, as shown in Table 6. Specifically, we remove each of the following strategies and measure the accuracy: Regularization (Section 4.1.2), Localization (Section 4.3), and Sample Selection (Section 4.2). The experiments demonstrate that selecting proper knowledge neurons to be updated is helpful in both handling interconnected knowledge and maintaining other knowledge. In addition, we reveal that Sample Selection significantly increases TA and SA, mitigating overfitting and shortcut unlearning issues.
Default 81.82 85.99 79.63 48.67 -
KLUE 36.70 82.97 74.69 58.16 69.78
(-) Regularization 40.40 79.74 67.59 51.24 64.54
(-) Localization 46.46 81.68 68.52 53.51 64.31
(-) Sample Selection 37.37 75.86 62.96 56.05 64.37
Table 6: Ablation studies for KLUE
@@ 6 Conclusion
Our research identifies the limitations of existing unlearning benchmarks, which have not explored the interconnectedness of knowledge. To overcome this issue, we define superficial unlearning and propose a new benchmark, FaithUn, for evaluating generalization, multi-hop knowledge unlearning, and shortcut unlearning. Using this benchmark, we empirically demonstrate that existing unlearning methods are vulnerable to superficial unlearning. Furthermore, we propose a novel knowledge-localized unlearning method, KLUE, which regularizes neuron attribution to identify and update only context-relevant neurons. We demonstrate that it outperforms existing unlearning methods, effectively mitigating superficial unlearning. Our paper first illuminates the phenomenon of superficial unlearning and raises a new research question for a deeper analysis of the unlearning field.
@@
FaithUn is constructed based on Wikidata and is designed to investigate the unlearning of knowledge about famous people for application in various language models. Although knowledge is more interconnected for well-known individuals, our benchmark does not examine a broader range of people. Our work does not evaluate knowledge editing methods, as knowledge editing and unlearning pursue different goals. Knowledge editing typically assesses models by post-edit accuracy, while unlearning emphasizes whether a model successfully forgets private or sensitive information. We therefore exclude existing knowledge editing methods from our experiments. Additionally, our study focuses solely on erasing the target label, leaving the issue of hallucinations in the unlearning process as future work, in line with prior studies.
@@
Acknowledgements
This work was supported by Adobe Research. This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea. The Institute of Engineering Research at Seoul National University provided research facilities for this work.
@@ A.1 Detailed Dataset Comparison
In this section, we present detailed comparisons with existing datasets to highlight the novelty of our benchmark. Our benchmark targets the unlearning of knowledge about well-known real-world entities, which are often memorized by language models, thereby addressing practical challenges in knowledge unlearning. Additionally, our benchmark captures the complex and interconnected nature of world knowledge by introducing three evaluation perspectives—generalization, multi-hop knowledge unlearning, and shortcut unlearning—for a more comprehensive analysis.
In summary, MUSE, KnowUnDo, and TOFU require fine-tuning to inject knowledge prior to unlearning, which limits their practicality. Furthermore, existing datasets—excluding RWKU and ours—fail to evaluate whether related knowledge to the target is appropriately preserved or removed during unlearning. However, RWKU also has limitations in that it only evaluates knowledge through varying textual expressions (e.g., cloze test and question answering) and related but semantically disjoint facts, thus overlooking deeper relational structures.
For example, RWKU includes a target sentence for unlearning: “Please forget Stephen King, who is an American author, renowned as the ‘King of Horror’”. It also presents a related question: “Who plays the character Jack Torrance in the film The Shining?”. RWKU evaluates whether an unlearned model preserves knowledge that is related but should not be removed after unlearning the target. In contrast, our benchmark introduces a more challenging setting by disentangling multiple pieces of knowledge about a single entity and evaluating whether the remaining knowledge is faithfully retained. We also assess the preservation of knowledge about other entities. These aspects are effectively evaluated by the TA and SA metrics. Moreover, RWKU has addressed only isolated facts with no direct knowledge dependency. By contrast, FaithUn is designed to evaluate unlearning in more realistic scenarios by incorporating multi-hop questions and handling directly connected pieces of knowledge.
A.2 Details in Dataset Construction
A.2.1 Selected Entities and Relations.
A.2.2 Dataset Analysis.
Dataset Format.
Our FaithUn benchmark includes four types of datasets: the Base QA dataset, the Paraphrased QA dataset, the Multi-hop QA dataset, and the Same-answer QA dataset. Each instance in the Base QA dataset is matched with instances in other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) to examine the impact of unlearning on these datasets. Dataset statistics for the FaithUn benchmark are shown in Table 7. Examples in the FaithUn benchmark are shown in Table 12.
@@ A.2 Details in Dataset Construction
Table 7: Dataset statistics. Each question in the Base QA dataset forms a cluster, and questions from other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) are mapped to those in the Base QA dataset, thereby being assigned to the corresponding cluster for evaluation.
@@
A.2.3 Question Generation Prompt Templates
@@ B.3 More Details for the Experiments
[…]
@@
(1) Gradient Ascent (GA): Unlike the gradient descent used during the pre-training phase, GA (Jang et al., 2023; Yao et al., 2023) maximize the negative log-likelihood loss on the forget set. This method helps shift the model away from its original predictions, aiding in the unlearning process. (2) Gradient Ascent with a Retaining Loss (GAret): GA tends to unlearn other unrelated knowledge since it just maximizes the negative log-likelihood loss on the forget set. Therefore, we add an auxiliary retention loss to maximize the log-likelihood of the retaining set, securing the retention of other irrelevant knowledge. (3) Direct Preference Optimization (DPO): We adopt preference optimization to unlearn a language model to generate another answer. DPO (Rafailov et al., 2024; Jin et al., 2024) utilizes positive and negative instances to train the model. Therefore, we select the correct answer as the negative instance and also define two types of DPO methods to determine positive ones: (1) DPOmis (DPO using a mismatched answer) and (2) DPOneg (DPO using a rejection answer). DPOmis utilizes a randomly sampled answer as the positive instance. On the other hand, DPOrej utilizes a rejection text “I can’t answer the question." as the positive instance. Two DPO methods both aim to increase the probability of the positive instance compared to the negative one for the forget set, and they switch the positive and negative instances for training the retaining set. We search for $\beta_{DPO}\in[0.1,0.5]$ to optimize models. (4) NPO: NPO is a modified version of DPO that exclusively retains negative examples without positive ones. NPO can also be explained as a straightforward modification of the GA loss. We implement NPO (Zhang et al., 2024) for extended experiments. We search for $\beta_{NPO}\in[0.1,0.5]$ to optimize models. (5) RMU: We implement RMU (Li et al., 2024), the representation learning-based unlearning model. For RMU experiments, we search for $\alpha_{RMU}\in\{20,50,100,150,200,300\}$ and use hyper-parameters $c=20$ and $l=7$, following the implementation details on the original GitHub Page [^3] https://github.com/centerforaisafety/wmdp. (6) Knowledge-Localized Unlearning (KLUE): We select only 5% of neurons from Feed-forward networks for the knowledge neuron localization, and update them using general gradient ascent with retention loss. We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term. The experiments analyzing varying hyper-parameters are shown in Section 5.4, Appendix C.2, and Appendix C.3.
Appendix C Additional Experiments
C.1 Sequential vs. Batch Unlearning
@@ Appendix C Additional Experiments
C.2 Hyper-parameter ($\alpha$) Experiments
@@
C.3 Neuron Ratio ($p$) Experiments
@@
C.4 Various Prompt Templates Experiments
@@
C.5 3-hop Questions Experiments
We conduct experiments on 3-hop questions to evaluate whether unlearning methods can erase the knowledge of 3-hop questions. Consequently, we reveal that every method does not effectively unlearn the 3-hop knowledge. It is because 3-hop questions inherently form an unnatural format which is not used in practical scenarios, such as: "What is the highest point on the continent where Barack Obama holds citizenship?".
[…]
Default 48.67 49.70
GA 48.34 48.32
GAret 53.21 49.09
RMU 53.05 47.49
KLUE 58.16 49.66
Table 10: Extended multi-hop (two and three hops) results for Gemma-2 (2B).
C.6 Extended Studies on Mismatched Pairs
We use mismatched pairs to adopt superficial knowledge regularization. For a deeper analysis of the regularization term, we conduct experiments on “When constructing mismatched QA pairs, why not use the same question but a randomly sampled answer?” as presented in the table below. We denote the new implementation using mismatched pairs of the same question but a random answer as KLUE (new). We observed that subtracting attribution scores based on the same question but a randomly sampled answer leads to degraded performance. This result suggests that when the original question is used in the knowledge regularization term, it inadvertently removes contextual knowledge rather than superficial knowledge, since it includes much of the contextual information.